OpenStudy (inkyvoyd):

Does the integration of sec^3 x start with the derivative of sec x tan x? I've heard that this integration problem is difficult, so I decided to try it out. Please just tell me if I'm on the right track; I don't need a walkthrough. THanks :D

OpenStudy (blockcolder):

Not sec x tan x, just sec x. :D

OpenStudy (inkyvoyd):

The derivative of sec x?

OpenStudy (blockcolder):

Yes. It's an integration by parts problem where u= sec(x).

OpenStudy (inkyvoyd):

but but but http://www.wolframalpha.com/input/?i=derive+sec+x+tan+x

OpenStudy (inkyvoyd):

I'ma be a scumbag and try it with the derivative of sec x tan x.

OpenStudy (blockcolder):

It's okay to do that if you know the antiderivative of sec(x).

OpenStudy (inkyvoyd):

Yes, I do.

OpenStudy (blockcolder):

Then it's a perfectly okay method. In fact, this is the first time I've seen that solution to finding the antiderivative of sec^3(x), so kudos to you! :D

OpenStudy (inkyvoyd):

Thanks, I came up with it (ironically enough) trying to find the antiderivative of sec x

OpenStudy (inkyvoyd):

(wanted to find the area between x=y and sqrt(x^2-1) and see if it was bounded as x approached infinity)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integrate+sec%5E3+x Click on "Show steps"

OpenStudy (inkyvoyd):

I will NEVA click on that link! :D

OpenStudy (blockcolder):

So you were figuring out whether \[\int_{1}^{\infty} x-\sqrt{x^2-1}\ \text{d}x\] converged, right? So you used x=sec(theta) and lo, a wild sec^3(x) appeared. =))

OpenStudy (inkyvoyd):

No, I was doing the stupid, "Let's look for the derivatives of random functions and hope they come out to sec x!". Of course it didn't work (didn't even come close; I didn't use ln()), but derived sec x to get sec x tan x, then sec x tan x to get sec x(sec^2 x+tan^2 x), and realized that tan^2 x=sec^2 x-1, so I got sec x(2 sec^2 x-1), or sec^3 x-sec x

OpenStudy (inkyvoyd):

*2 sec^3 x- sec x

OpenStudy (inkyvoyd):

It's amazingly easy to use that to integrate sec^3 x; unfortunately, I did not know how to integrate sec x LOL

OpenStudy (inkyvoyd):

I eventually asked my sister, who had forgot, but later remembered to set u=(sec x+tan x)du

OpenStudy (inkyvoyd):

then, the usual integral du/u, or ln|u|, which resubstituted into ln|sec x+tan x|

OpenStudy (inkyvoyd):

@blockcolder

OpenStudy (blockcolder):

Yeah, and you'd get the formula for integral of sec^3(x). =))

OpenStudy (inkyvoyd):

I was amazed by the discovery actually; that a little bit of playing around gives you the answer to a "very hard" integration problem.

OpenStudy (inkyvoyd):

Anyone else got easy ways to do this integration method?

OpenStudy (anonymous):

look in the back page of your book for the "trig reduction formula" and you will find it